# thanks: Corresponding author

### v1

@@ II.2 Mutually equi-biased bases for $d=2$
$\alpha=\cos^{-1}[\delta/(1+\delta)].$ (6)
Here, $\delta=q_{0}-q_{1}$ is introduced both as an independent parameter for the binary distribution $q=(q_{0},q_{1})$ (see Appendix C for the underlying motivation) and as a measure of its unbiasedness.
@@
Accordingly, the Bloch vectors form a symmetric or equiangular triple such that they are mutually symmetric with the same mutual angle $\delta$ (see Fig. 1).
Figure 1: Bloch vector representation for MEB, when $d=2$. Geometrically, equi-biasedness means equal angle between Bloch vectors of different bases.
@@ II.3 Mutually equi-biased bases for $d=3$
Mutual orthogonality of the rows (and columns) gives simply the following conditions
@@
A subset of the triangle probability that satisfies the chain-link conditions (13)—so that the matrix $\mathcal{Q}$ is unistochastic and can be generated from a unitary matrix—is shown in Fig. 2(a). The figure is plotted in the $\delta_{2}\delta_{1}$-plane, where the two independent parameters $\delta_{1}$ and $\delta_{2}$ are defined as follows (see Appendix C)
$q_{0}$ $=$ $1/3+2\delta_{1}/3,$ (17)
$q_{1}$ $=$ $1/3-\delta_{1}/3-\delta_{2}/\sqrt{3},$
$q_{2}$ $=$ $1/3-\delta_{1}/3+\delta_{2}/\sqrt{3}.$
Equivalently
$\delta_{1}=(3q_{0}-1)/2,\qquad\delta_{2}=\sqrt{3}(q_{2}-q_{1})/2,$ (18)
such that
$-1/2\leq$ $\delta_{1}$ $\leq 1,$ (19)
$(-1+\delta_{1})/\sqrt{3}\leq$ $\delta_{2}$ $\leq(1-\delta_{1})/\sqrt{3}.$ (20)
Figure 2: Geometric representation of the probability distribution $q=(q_{0},q_{1},q_{2})$ in the $\delta_{2}\delta_{1}$-plane. (a) A subset of the triangle probability where the bistochastic matrix $\mathcal{Q}$ is a unistochastic, i.e., the probabilities for which the conditions (13) are satisfied. (b) The feasible region that satisfy the constraint $\mu\leq 5/9$. (c) The feasible domain, after imposing the chain-linked conditions (26). (d) The portion of the feasible region for which $\mu\leq 1/2$.
One can construct the second unitary matrix $\mathcal{V}$ with the same prescribed moduli of its matrix elements as $\mathcal{U}$ as $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$. Regarding the probability $q=(q_{0},q_{1},q_{2})$, although both $\mathcal{U}$ and $\mathcal{V}$ generate bases that are equi-biased with respect to the standard basis, they are not necessarily equi-biased with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same probability matrix $\mathcal{Q}$, up to a permutation of its entries. This places some constraints on phases $\alpha_{1}$ and $\alpha_{2}$. Furthermore, the mutual equi-biasedness imposes some constraints on the index $\mu=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}$ of the distribution $q=(q_{0},q_{1},q_{2})$—Not any value of $\mu$ is allowed to generate a bistochastic matrix from unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$.
Here, $\sigma\in S_{3}$ denotes a permutation of three elements $\{0,1,2\}$, where $S_{3}$ is the symmetric group of degree $3$. For further use, recall that the $3!$ group operations of $S_{3}$ can be denoted as: the identity $\mathrm{id}$, the pair permutations $(12)$, $(13)$, $(23)$, and the forward and backward cycles $(123)$ and $(321)$. After some calculations, we get
$\cos{\alpha_{1}}+\cos{\alpha_{2}}+\cos{\alpha_{12}}=\frac{1-3\mu}{1-\mu}.$ (25)
For $\mu\in[1/3,1)$, the rhs of Eq. (25) can take any nonpositive value $(-\infty,0]$, but the lhs is restricted to the interval [3/2,3][-3/2,3]. This poses restrictions on the probability distribution $q=(q_{0},q_{1},q_{2})$ in the sense that from the set of all distributions that can generate unistochastic matrices, i.e., those that satisfy the triangle inequalities (13), only distributions with index $\mu\in[1/3,5/9]$ can generate unistochastic matrices from $\mathcal{V}^{\dagger}\mathcal{U}$. The resulting feasible region is depicted in the Fig. 2(b).
Note, however, that although the index condition $\mu\leq 5/9$ is necessarily to generate a bistochastic matrix from the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$, it is not sufficient. Indeed, the conditions (22) to (24) pose some chain-link conditions on four segments $q_{0},q_{1},q_{2}$, and $\sqrt{q_{i}}$, i.e., the following six nontrivial conditions hold
$q_{i}\leq q_{j}+q_{k}+\sqrt{q_{j}},$ (26)
for $(i,j,k)$ taken from $(0,1,2)$. After imposing these constraints, the resulting feasible domain is reduced to the region shown in the Fig. 2(c). For the reason become clear soon, we have to know the region in which $\mu\leq 1/2$. Figure 2(d) illustrates the portion of the feasible region for which $\mu\in[1/3,1/2]$. We are now in a position to characterize a complete set of MEBs for $d=3$.
@@
Table 1 presents all possible choices for the distribution $q=(q_{0},q_{1},q_{2})$ (first column) versus all possible choices for two pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ (first row). For each chosen probability and each chosen pairs of phases, three equivalent distributions for which three unitary matrices $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ generate equivalent unistochastic matrices are provided. Remarkably, for a given distribution $q=(q_{0},q_{1},q_{2})$, two other distributions corresponding to cyclic permutations $(123)$ and $(321)$, i.e., $q^{(123)}=(q_{1},q_{2},q_{0})$ and $q^{(321)}=(q_{2},q_{0},q_{1})$, have also contribution to generate equivalent unistochastic matrices $\mathcal{Q}$, $\mathcal{Q}^{(123)}$, and $\mathcal{Q}^{(321)}$.
@@
II.3.1 Proof of Proposition 2
Proof.
Equation (25) imposes a constraint on the allowed values of the phases $\alpha_{1}$ and $\alpha_{2}$ under which two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ are mutually equi-biased. It depends only on $\mu$, but not on where the probability $q=(q_{0},q_{1},q_{2})$ is located in the feasible region. This makes the constraint invariant under the symmetric group $S_{3}$. A contour plot of this equation in $\alpha_{1}\alpha_{2}$-plane for different values of $\mu$ is illustrated in Fig. 3. The contours which are bounded within the hexagon correspond to $\mu\in[1/3,1/2]$, and those outside the hexagon, i.e., within the two outer triangles, correspond to $\mu\in(1/2,5/9]$.
The feasible region defined above for the probability distribution $q=(q_{0},q_{1},q_{2})$ does not guarantee solutions at any point of the feasible region. Indeed, symmetry conditions requires that the solutions exist only within a measure zero subset of the feasible region, defined by six curves. To see this, note that although the constraint (25) possesses the full symmetry $S_{3}$, a look at Eqs. (22) to (24) reveals that the LHSs of the set is invariant under a subgroup of $S_{3}$, namely, the cyclic group $C_{3}=\{\mathrm{id},(123),(321)\}$. The set, however, will possess the full symmetry under $S_{3}$ if and only if two out of the three terms $\cos{\alpha_{1}}$, $\cos{\alpha_{2}}$, and $\cos{\alpha_{12}}$ become equal. This results in three possible cases as
  1. (a) :
    $\alpha_{1}=0,\quad\quad\alpha_{2}=\theta^{\mu}$,
  2. (b) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=0$,
  3. (c) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=\theta^{\mu}$,
Similarly, for cases (b) and (c), the set of Eqs. (22) to (24) reduce to
and
respectively. In each case, the corresponding set of equations gives, up to a permutation, the same set of solutions (29) and (30) for the probabilities $(q_{0},q_{1},q_{2})$. Irrespective of which case (a), (b), or (c) is considered, only three out of six possible permutations $\sigma$, namely, those corresponding to the cyclic group $C_{3}=\{\mathrm{id},(123),(321)\}$, are contributed in the solutions.
To complete the characterization of the set of MEBs, we have to construct the third unitary matrix $\mathcal{W}$ as $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$ where $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$. Similar to the arguments given above, we require that the unitaries $\mathcal{W}\mathcal{U}$ and $\mathcal{W}\mathcal{V}$ generate the unistochastic matrix $\mathcal{Q}$, up to a permutation of its rows and columns. The first requirement is achieved if we select the pair $(\beta_{1},\beta_{2})$ from the cases (a), (b), or (c), mentioned above for $(\alpha_{1},\alpha_{2})$. To ensure that $\mathcal{W}\mathcal{V}$ also generates the unistochastic matrix $\mathcal{Q}$, we have to choose the pairs $(\beta_{1},\beta_{2})$ distinct from the pairs $(\alpha_{1},\alpha_{2})$. All possible solutions are listed in Table 1.
Figure 4: The curves represent the allowed values of the distribution $q=(q_{0}=q^{\mu}_{+}$, $q_{1}=\frac{1-\mu}{2}$, $q_{2}=q^{\mu}_{-})$, or any permutation of its probabilities, when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$. Each curve starts from $\mu=1/3$ at the center of the feasible region and ends at $\mu=1/2$ on the boundary of the feasible region. The circles correspond to regions with constant index $\mu$. Corresponding to each $\mu$, only six points, i.e., the distribution $q=(q_{0}=q^{\mu}_{+}$, $q_{1}=\frac{1-\mu}{2}$, $q_{2}=q^{\mu}_{-})$ and its permutations, are allowed for a given $\mu$.
II.3.2 Example: $d=3$ and $\mu=1/2$

### v2

@@ II.2 Mutually equi-biased bases for $d=2$
$\alpha=\cos^{-1}[\delta/(1+\delta)],$ (16)
where the unbiasedness parameter $\delta=q_{0}-q_{1}$ is defined by Eq. (7).
@@
Accordingly, the Bloch vectors form a symmetric or equiangular triple such that they are mutually symmetric with the same mutual angle $\vartheta=\cos^{-1}[\delta]$ (see Fig. 1).
Figure 1: Bloch vector representation for MEB, when $d=2$. Geometrically, equi-biasedness means equal angle, $\vartheta=\cos^{-1}[\delta]$, between Bloch vectors of different bases.
@@ II.3 Mutually equi-biased bases for $d=3$
The unitarity condition $\mathcal{U}\mathcal{U}^{\dagger}={\mathbb{I}}$, i.e., the mutual orthogonality of the rows (and columns), gives simply the following conditions
@@
For each probability distribution $q=(q_{0},q_{1},q_{2})$ satisfying the chain-link conditions (23), the associated unitary matrix $\mathcal{U}$, defined in Eq. (18) with phases specified in Eqs. (24) to (26), generates the corresponding unistochastic matrix $\mathcal{Q}$. By construction, the matrix $\mathcal{U}$ represents the most general form consistent with $q=(q_{0},q_{1},q_{2})$ and defines the second orthonormal basis $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ with respect to the standard basis $\{|e_{k}^{(1)}\rangle\}$. Figure 2(a) shows probability triangle with its unistochastic subset. The boundaries of this subset are hypocycloids, each corresponding to the saturation of one of the chain-link inequalities (23). A complete clarification of Fig. 2, along with the additional conditions leading to the feasible region, is provided in Remark 3.
Given $\mathcal{U}$, we have to construct two unitary matrices, $\mathcal{V}$ and $\mathcal{W}$, such that all bases are mutually equi-biased. The following proposition provides a complete set of four MEBs for $d=3$.
Figure 2: Geometric representation of the probability distribution $q=(q_{0},q_{1},q_{2})$ and the construction of the feasible region. (a) Equilateral triangle associated to probability distribution $q=(q_{0},q_{1},q_{2})$, with its unistochastic subset. All subfigures are plotted in the $\delta_{2}\delta_{1}$-plane. Each subfigure corresponds to a constraint applied in succession. See Remark 3 for details.
@@
Before proceeding with the proof, we illustrate the form of the solutions in Fig. 3. They represent a one-parameter class of probability distributions $q=(q_{0},q_{1},q_{2})$ within the feasible region of the probability triangle, defined by six curves each starting at $\mu=1/3$ at the center of the feasible region and ending at $\mu=1/2$ on the boundary. The curves represent the allowed values of the distribution $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$, or any permutation of its probabilities, for $(\alpha_{1},\alpha_{2})$ chosen from the pairs given in Eq. (27).
Figure 3: Geometrical representation of the probability distribution $q=(q_{0},q_{1},q_{2})$, for which a complete set of four MEBs exist. The figure is plotted in the $\delta_{2}\delta_{1}$-plane, where independent parameters $\delta_{1}$ and $\delta_{2}$ are defined in Eq. (8). The curves represent the allowed values of the distribution $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$, or any permutation of its probabilities, when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$. The circles correspond to regions with constant index $\mu$. Corresponding to each $\mu$, only six points, i.e., the distribution $q$ and its permutations, are allowed for a given $\mu$.
Proof.
Having the first unitary matrix $\mathcal{U}$, we have to find two other unitary matrices, $\mathcal{V}$ and $\mathcal{W}$, with the same moduli of matrix elements as $\mathcal{U}$. The second unitary matrix $\mathcal{V}$ can be constructed as $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$. Regarding the probability $q=(q_{0},q_{1},q_{2})$, although both $\mathcal{U}$ and $\mathcal{V}$ generate bases that are equi-biased with respect to the standard basis, they are not necessarily equi-biased with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same probability matrix $\mathcal{Q}$, up to a permutation of its entries. This places some constraints on phases $\alpha_{1}$ and $\alpha_{2}$.
Here, $\sigma\in S_{3}$ denotes a permutation of three elements $\{0,1,2\}$, where $S_{3}$ is the symmetric group of degree $3$. For clarity, the six operations of $S_{3}$ can be written as: the identity $\mathrm{id}$, the transpositions $(01)$, $(02)$, $(12)$, and the forward and backward cycles $(012)$ and $(210)$. After some calculations, we get
$\cos{\alpha_{1}}+\cos{\alpha_{2}}+\cos{\alpha_{12}}=\frac{1-3\mu}{1-\mu},$ (35)
which holds for $\mu\in[1/3,5/9]$.
Equation (35) imposes a constraint on the allowed values of the phases $\alpha_{1}$ and $\alpha_{2}$ under which two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ are mutually equi-biased. A contour plot of this equation in $\alpha_{1}\alpha_{2}$-plane for different values of $\mu$ is illustrated in Fig. 4. The contours which are bounded within the hexagon correspond to $\mu\in[1/3,1/2]$, and those outside the hexagon, i.e., within the two outer triangles, correspond to $\mu\in(1/2,5/9]$.
Clearly, the constraint imposed by Eq. (35) is invariant under the symmetric group $S_{3}$; it depends only on $\mu$ but not on where the probability $q=(q_{0},q_{1},q_{2})$ is located in the feasible region. However, a look at the LHSs of Eqs. (32) to (34) reveals that this set is invariant under a subgroup of $S_{3}$, namely, the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$. The set, however, will possess the full symmetry under $S_{3}$ if and only if two out of the three terms $\cos{\alpha_{1}}$, $\cos{\alpha_{2}}$, and $\cos{\alpha_{12}}$ become equal. This results in three possible cases as
  1. (a) :
    $\alpha_{1}=0,\quad\quad\alpha_{2}=\theta^{\mu}$,
  2. (b) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=0$,
  3. (c) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=\theta^{\mu}$,
Depending on the choice of the permutation $\sigma\in S_{3}$, the set of equations determines the probability distribution $q=(q_{0},q_{1},q_{2})$. More precisely, the choice $\sigma=\mathrm{id}$ fixes $q_{1}=q^{\mu}_{0}$, so that $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$, where $q^{\mu}_{0}=(1-\mu)/2$ and $q^{\mu}_{\pm}$ is given in Eq. (30). Similarly, the cycle $\sigma=(012)$ fixes $q_{0}=q^{\mu}_{0}$, giving $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$, while the cycle $\sigma=(210)$ yields $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$. What about the transpositions $(01)$, $(02)$, and $(12)$? In these cases, the set possesses solution only for $\mu=1/3$, corresponding to the uniform distribution $q=(1/3,1/3,1/3)$.
In a similar manner, for cases (b) and (c), the set of Eqs. (32) to (34) reduce to
and
respectively. In each case, the corresponding set of equations gives, up to a permutation, the same set of solutions (29) and (30) for the probabilities $(q_{0},q_{1},q_{2})$. Irrespective of which case (a), (b), or (c) is considered, only three out of six possible permutations $\sigma$, namely, those corresponding to the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$, are contributed in the solutions. Table 1 presents all possible solutions of Eqs. (32) to (34), corresponding to three cases (a), (b), (c), and three permutations $\mathrm{id}$, $(012)$, $(210)$.
Table 1: Possible solutions of Eqs. (32) to (34).
$\sigma=\mathrm{id}$ $\sigma=(012)$ $\sigma=(210)$
$\alpha=(0,\theta^{\mu})$ $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$ $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$ $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$
$\alpha=(\theta^{\mu},0)$ $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$ $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$ $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$
$\alpha=(\theta^{\mu},\theta^{\mu})$ $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$ $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$ $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$
To complete the characterization of the set of MEBs, we have to construct the third unitary matrix $\mathcal{W}$ as $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$ where $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$. Similar to the arguments given above, we require that the unitaries $\mathcal{W}\mathcal{U}$ and $\mathcal{W}\mathcal{V}$ generate the unistochastic matrix $\mathcal{Q}$, up to a permutation of its rows and columns. The first requirement is achieved if we select the pair $(\beta_{1},\beta_{2})$ from the cases (a), (b), or (c), mentioned above for $(\alpha_{1},\alpha_{2})$. To ensure that $\mathcal{W}\mathcal{V}$ also generates the unistochastic matrix $\mathcal{Q}$, we have to choose the pairs $(\beta_{1},\beta_{2})$ distinct from the pairs $(\alpha_{1},\alpha_{2})$. ∎
It is clear from the construction that, although all three unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate the same unistochastic matrix $Q$, associated with the probability distribution $q=(q_{0},q_{1},q_{2})$, the unistochastic matrices resulting from their mutual products $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ are not necessary equal to $Q$. However, they are equivalent in the sense that they are generated either by $q=(q_{0},q_{1},q_{2})$ or by a permutation of its components, i.e., $q^{\sigma}=(q_{\sigma(0)},q_{\sigma(1)},q_{\sigma(2)})$, such as $q^{(012)}=(q_{1},q_{2},q_{0})$ and $q^{(210)}=(q_{2},q_{1},q_{0})$. All such possibilities are summarized in Table 2. The first column represents all possible choices for the distribution $q=(q_{0},q_{1},q_{2})$, and the first row represents all possible choices for two pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$. For each chosen probability and each chosen pairs of phases, three equivalent distributions for which three unitary matrices $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ generate equivalent unistochastic matrices are provided.
@@
As we have seen, the mutual equi-biasedness imposes some constraints on the index $\mu=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}$ of the distribution $q=(q_{0},q_{1},q_{2})$, meaning that, not any value of $\mu$ is allowed to generate a bistochastic matrix from the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$. The following remark clarifies how the feasible region is derived.
Remark 3.
The feasible region within the probability triangle is determined by applying a sequence of constraints arising from physical considerations. These conditions successively restrict the parameter space, resulting in a subset that satisfies all necessary properties. Figure 2 outlines how the feasible region is constructed in this case, with each subfigure depicting a specific constraint imposed on the probability triangle, as detailed below.
Figure 2(a): A subset of the probability triangle that satisfies the chain-link conditions (23) is shown in Fig. 2(a). Corresponding to each probability distribution $q=(q_{0},q_{1},q_{2})$ within the subset, the associated unitary matrix $\mathcal{U}$ generates the corresponding unistochastic matrix $\mathcal{Q}$.
Figure 2(b): For $\mu\in[1/3,1)$, the rhs of Eq. (35) can take any nonpositive value $(-\infty,0]$, but the lhs is restricted to the interval [3/2,3][-3/2,3]. This poses restrictions on the index $\mu$ in the sense that only distributions with index $\mu\in[1/3,5/9]$ can generate unistochastic matrices from $\mathcal{V}^{\dagger}\mathcal{U}$. The resulting feasible region is depicted in the Fig. 2(b).
Figure 2(c): Equations (32) to (34) pose some chain-link conditions on the four segments $q_{0},q_{1},q_{2}$, and $\sqrt{q_{i}}$ as well, i.e., the nontrivial conditions $q_{i}\leq q_{j}+q_{k}+\sqrt{q_{j}}$ hold, for $(i,j,k)$ taken from $(0,1,2)$. After imposing these constraints, the resulting feasible domain is reduced to the region shown in the Fig. 2(c).
Figure 2(d): Finally, Eq. (28), together with the condition $|\cos{\theta^{\mu}}|\leq 1$, leads to the constraint $\mu\leq 1/2$. Figure 2(d) illustrates the feasible region corresponding to $\mu\in[1/3,1/2]$, consistent with all previously imposed constraints.
II.3.1 Example: $d=3$ and $\mu=1/2$